# GUT-Scale Smooth Hybrid Inflation with a Stabilized Modulus in Light of ACT and SPT Data

### v1

@@ 5.3 Monotonicity of $V_{\rm I}$
An outstanding feature of our proposal is that $V_{\rm I}$ in Eq. (34) remains monotonic during smFHI for both SUGRA settings. To highlight it we display the variation [the derivative w.r.t $\sigma$] of $V_{\rm I}$ as a function of $\sigma$ in the left [right] plot of Fig. 4. We focus on smFHI-I with $p=2$. For shSUGRA (solid lines) we use the parameters shown in the column with $p=2$ in Table 1 whereas for NSUGRA (dashed lines) we use $\alpha=1/15$ and $\beta=-1$ resulting to $N_{0}\simeq-1.76$ and $n_{\rm s}=0.9723$. The values of $\sigma_{\star}$ and $\sigma_{\rm f}$ are also depicted. Note that $\sigma_{\star}$ and $\sigma_{\rm f}$ coincide for the two presented lines and so our assumption for the derivation of Eq. (56) is justified. Most importantly, it is clear that $V_{\rm I}$ remains a monotonically increasing function of $\sigma$ in both cases since $V_{\rm I,\sigma}>0$. So, unnatural restrictions on the initial conditions for smFHI due to the appearance of a maximum and a minimum of $V_{\rm I}$ can be avoided.
@@ 6 Conclusions
Let us, finally, note that a complete inflationary scenario must also account for the transition to a radiation-dominated universe and the generation of the observed baryon asymmetry. These aspects have been extensively explored in the context of smFHI with canonical or quasi-canonical Kähler potentials – see, e.g., Refs. [5, 9, 15, 10]. Our setup preserves many of the successful features of this post-inflationary evolution, which may impose additional constraints on the viable parameter space and help identify the most compelling version of smFHI.
@@ Appendix A Stabilization of the Modulus Before smFHI
$V_{\rm FI}=g^{2}m_{\rm P}^{4}\left(q_{h}\beta\ln|h/m_{\rm P}|^{-2}+m_{\rm FI}^{2}/m_{\rm P}^{2}\right)^{2},$ (61)

### v2

@@ 5.3 Monotonicity of $V_{\rm I}$
An outstanding feature of our proposal is that $V_{\rm I}$ in Eq. (34) remains monotonic during smFHI for both SUGRA settings. To highlight it, we display the variation [the derivative w.r.t $\sigma$] of $V_{\rm I}$ as a function of $\sigma$ in the left [right] plot of Fig. 4. We focus on smFHI-I with $p=2$. For shSUGRA (solid lines) we use the parameters shown in the column with $p=2$ in Table 1 whereas for NSUGRA (dashed lines) we use $\alpha=1/15$ and $\beta=-1$ resulting to $N_{0}\simeq-1.76$ and $n_{\rm s}=0.9723$. The values of $\sigma_{\star}$ and $\sigma_{\rm f}$ are also depicted. Note that $\sigma_{\star}$ and $\sigma_{\rm f}$ coincide for the two presented lines and so our assumption for the derivation of Eq. (57) is justified. Most importantly, it is clear that $V_{\rm I}$ remains a monotonically increasing function of $\sigma$ in both cases since $V_{\rm I,\sigma}>0$. So, unnatural restrictions on the initial conditions for smFHI due to the appearance of a maximum and a minimum of $V_{\rm I}$ can be avoided.
@@ 6 Conclusions
Let us, finally, note that a complete inflationary scenario must also account for the transition to a radiation-dominated universe and the generation of the observed baryon asymmetry. In Appendix C we provide a possible path to this direction which shows that our setup preserves many of the successful features of the post-inflationary evolution extensively explored in the context of smFHI with other Kähler potentials – see, e.g., Refs. [5, 9, 15, 10]. On the other hand, this stage of the cosmological history may impose additional constraints on the viable parameter space and helps identify the most compelling version of smFHI.
@@ Appendix A Stabilization of the Modulus Before smFHI
$V_{\rm FI}=g^{2}m_{\rm P}^{4}\left(q_{h}\beta\ln|h/m_{\rm P}|^{-2}+m_{\rm FI}^{2}/m_{\rm P}^{2}\right)^{2}/2,$ (62)
@@ Appendix B GUT-Scale smFHI within mSUGRA
Appendix C Reheating Process, Lepton-Asymmetry and Gravitino Abundances
We exemplify here a possible post-inflationary completion of our models which clarifies the fact that $S$ oscillations dominate over those of the Higgs fields ($\bar{\Phi}-\Phi$ or $\Psi$) and verifies that a scenario of non-thermal leptogenesis consistent with the gravitino ($\widetilde{G}$) constraint is implementable.
We focus on smFHI-I which allows for a direct coupling of $\bar{\Phi}$ to right-handed neutrinos $N^{c}_{i}$ via the superpotential term [5]
$\lambda_{[N^{c}}i](\bar{\Phi}N^{c}_{i})^{2}/M_{\rm S}$ (67)
where $\bar{\Phi}$ represents the SM singlet direction in $\bar{\Phi}$. This coupling is compatible with the symmetries of $W$ if $2R(N^{c}_{i})=R(W)$ and $(B-L)(\bar{\Phi})=-(B-L)(N^{c}_{i})$. The coupling above provides $N^{c}_{i}$ with mass
$M_{[N^{c}}1]=2\lambda_{[N^{c}}1]\langle{\Phi}\rangle^{2}/M_{\rm S}$ (68)
and allows for the decay into $N^{c}_{i}$ of the inflaton system – which consists of $S$ and $\delta\Phi_{+}=(\delta\bar{\Phi}+\delta\Phi)/\sqrt{2}$, where $\delta\bar{\Phi}=\bar{\Phi}-\langle{\bar{\Phi}}\rangle$ and $\delta\Phi=\Phi-\langle{\Phi}\rangle$ – via the common decay width
$\Gamma_{\rm I}=\frac{1}{32\pi}\gamma^{2}m_{\rm I}~~\mbox{with}~~\gamma=M_{[N^{c}}1]/\sqrt{2}\langle{\Phi}\rangle.$ (69)
Here we assume that only the decay into the lightest $N^{c}_{i}$ is possible – without competitive decay channels into the MSSM Higgs superfields $H_{u}$ and $H_{d}$ as in Ref. [15, 71] – and
$10T_{\rm rh}\lesssim M_{[N^{c}}1]\lesssim m_{\rm I}/2.$ (70)
The out-of-equilibrium decay of $N^{c}_{i}$ generates an $L$-asymmetry, which is not erased due to the left-most inequality in Eq. (70), and obtains a maximal value [93]
$\varepsilon_{L}=-\frac{3}{8\pi}\frac{m_{\nu_{\tau}}M_{[N^{c}}1]}{\langle{H_{u}}\rangle^{2}}\>\>\mbox{where}\>\>m_{\nu_{\tau}}=\sqrt{\Delta m^{2}_{\oplus}}\simeq 0.05\>{\rm eV}$ (71)
is the mass of heaviest light neutrino $\nu_{\tau}$ and $\Delta m^{2}_{\oplus}$ [94] the atmospheric neutrino mass-squared difference. Also, we set $\langle{H_{u}}\rangle=174~{\mbox{\rm GeV}}$ adopting the large $\tan\beta$ regime of MSSM.
During the period of oscillations – when $m_{\rm I}\gg H_{\rm I0}$ with $H_{\rm I0}$ being defined in Eq. (36) – the energy densities $\rho_{S}$ and $\rho_{\delta\Phi_{+}}$ of the $S$ and $\delta\Phi_{+}$ respectively, the energy density $\rho_{\rm R}$ of the produced radiation, and the number densities $n_{L}$ of the leptons and $n_{\widetilde{G}}$ of the $\widetilde{G}$’s satisfy the following Boltzmann equations – cf. Ref. [71]:
$\dot{\rho}_{S}+3H\rho_{S}+(1+w_{S})\Gamma_{\rm I}\rho_{S}=0,$ (72a)
$\dot{\rho}_{\delta\Phi_{+}}+3H\rho_{\delta\Phi_{+}}+(1+w_{\delta\Phi_{+}})\Gamma_{\rm I}\rho_{\delta\Phi_{+}}=0,$ (72b)
$\dot{\rho}_{\rm R}+4H\rho_{\rm R}-(1+w_{S})\Gamma_{\rm I}\rho_{S}-(1+w_{\delta\Phi_{+}})\Gamma_{\rm I}\rho_{\delta\Phi_{+}}=0,$ (72c)
$\dot{n}_{L}+3Hn_{L}-2\varepsilon_{L}\Gamma_{\rm I}\left((1+w_{S})\rho_{S}+(1+w_{\delta\Phi_{+}})\rho_{\delta\Phi_{+}}\right)/m_{\rm I}=0,$ (72d)
$\dot{n}_{\widetilde{G}}+3Hn_{\widetilde{G}}-C_{\widetilde{G}}\left(n^{\rm eq}\right)^{2}=0.$ (72e)
Here the overdot denotes derivation with respect to the cosmic time $t$ and $H$ is the Hubble parameter during this period which is given by
$H=\frac{1}{\sqrt{3}m_{\rm P}}\left(\rho_{S}+\rho_{\delta\Phi_{+}}+\rho_{\rm R}\right)^{1/2}.$ (73)
The barotropic indices $w_{S}$ and $w_{\delta\Phi_{+}}$ can be approximately estimated from Eq. (41) taking into account Eq. (23) and setting $\sigma$ and $\phi$ close to their v.e.vs in Eq. (24). Namely, we find $n=2$ and so $w_{S}=0$ for $\varphi=\sigma$ in Eq. (41). Similarly, for $\varphi=\phi$ Eq. (23) yields $n=4p$ and so $w_{\delta\Phi_{+}}=(2p-1)/(2p+1)$. Also $C_{\widetilde{G}}$ and $n^{\rm eq}$ are a collision term for $\widetilde{G}$ production and the equilibrium number density of each bosonic relativistic species. They have respectively the form
$C_{\widetilde{G}}\simeq\frac{3\pi}{16\zeta(3)m_{\rm P}^{2}}\sum_{i=1}^{3}c_{i}g_{i}^{2}\ln\left(\frac{k_{i}}{g_{i}}\right)~~\mbox{and}~~n^{\rm eq}={\zeta(3)\over\pi^{2}}T^{3},$ (74)
where in writing $C_{\widetilde{G}}$ we take the limit of massless MSSM gauginos and of zero top trilinear coupling constant. We take [92] $(c_{i})=(21.336,35.33,29.4)$, $g_{i}$ are the gauge coupling constants of the MSSM, and $(k_{i})=(1.097,1.38,3.07)$. Finally, the temperature $T$ and the entropy density $\sf s$} are found from the relations
$\rho_{\rm R}=\frac{\pi^{2}}{30}g_{\rm rh*}T^{4}~~\mbox{and}~~{\sf s}=\frac{2\pi^{2}}{45}g_{\rm rh*}T^{3}.$ (75)
The system of Eqs. (72a)-(72e) can be solved under the following initial conditions
$\rho_{S}(0)=\rho_{\delta\Phi_{+}}(0)=V_{\rm I0}/2~~\mbox{and}~~\rho_{\rm R}(0)=n_{\widetilde{G}}(0)=n_{L}(0)=0,$ (76)
where we assumed that the inflationary energy density is equally distributed between the oscillatory subsystems $S$ and $\delta\Phi_{+}$.
[…]
From Fig. 5a we observe that smFHI is followed by the matter-dominated era driven by the decaying $S$ since the decay of $\delta\Phi_{+}$ occurs very early at $T=T_{\delta\Phi_{+}}\simeq 2\cdot 10^{10}~{\mbox{\rm GeV}}$ – this temperature corresponds to the intersection of the $\rho_{\delta\Phi_{+}}$ and $\rho_{\rm R}$ lines – due to the strong hierarchy between $w_{S}=0$ and $w_{\delta\Phi_{+}}=3/5$. The $S$ oscillations continue until $\rho_{S}$ meets $\rho_{\rm R}$ at $T_{\rm rh}=5.7\cdot 10^{8}~{\mbox{\rm GeV}}$. This numerical result is in excellent agreement with the estimate obtained by using $w_{\rm rh}=w_{S}$ in the approximate formula of $T_{\rm rh}$ in Ref. [71]. After reheating, the universe enters a conventional radiation dominated era. Therefore, although our scenario involves two oscillatory systems, $\delta\Phi_{+}$ and $S$, the final $T_{\rm rh}$ can be accurately computed by considering solely the system with the lowest $w$. Obviously, this conclusion remains valid for larger $p$ values too.
In Fig. 5b we depict the cosmological evolution of the absolute values of the lepton-asymmetry and $\widetilde{G}$ abundances defined as follows
$Y_{L}=n_{L}/{\sf s}~~\mbox{and}~~Y_{\widetilde{G}}=n_{{\widetilde{G}}}/{\sf s}.$ (77)
We see that $|Y_{L}|$ and $Y_{\widetilde{G}}$ reach their constant values equal to $2.3\times 10^{-10}$ and $1.8\times 10^{-13}$ respectively immediately after the completion of $S$ decay, close to $T\simeq T_{\rm rh}$ – note that $Y_{L}$ originates almost exclusively from the lepton asymmetry $2\varepsilon_{L}$ generated by the $S$ decay. $Y_{L}$ is partially converted through sphaleron effects into a yield of the $B$ asymmetry of the universe $Y_{B}$ which meets its observational value [28] since $0.35Y_{L}(T_{\rm rh})=8.7\cdot 10^{-11}$. Moreover, the present value $Y_{\widetilde{G}}$ is compatible with the upper bound posed by the requirements [95] of Big Bang nucleosynthesis for $m_{\widetilde{G}}\gtrsim 10~{\mbox{\rm TeV}}$ – cf. Ref. [71]. Therefore, non-thermal leptogenesis is a realistic possibility within smFHI-I.
The case of smFHI-II requires further investigation since the coupling in Eq. (67) has to be modified and so the reheating process, the generation of neutrino masses and the leptogenesis mechanism depend on the structure of the complete model – see, e.g., Ref. [9, 10].